# Multimodal Learning with Uncertainty Quantification based on Discounted Belief Fusion

### v1

@@ Abstract
Multimodal AI models are increasingly used in fields like healthcare, finance, and autonomous driving, where information is drawn from multiple sources or modalities such as images, texts, audios, videos. However, effectively managing uncertainty—arising from noise, insufficient evidence, or conflicts between modalities—is crucial for reliable decision-making. Current uncertainty-aware ML methods leveraging, for example, evidence averaging, or evidence accumulation underestimate uncertainties in high-conflict scenarios. Moreover, the state-of-the-art evidence averaging strategy struggles with non-associativity and fails to scale to multiple modalities. To address these challenges, we propose a novel multimodal learning method with order-invariant evidence fusion and introduce a conflict-based discounting mechanism that reallocates uncertain mass when unreliable modalities are detected. We provide both theoretical analysis and experimental validation, demonstrating that unlike the previous work, the proposed approach effectively distinguishes between conflicting and non-conflicting samples based on the provided uncertainty estimates, and outperforms the previous models in uncertainty-based conflict detection.
@@ 1 Introduction
Recently, the use of Artificial Intelligence (AI) has surged dramatically for both autonomous and assisted decision-making. In fields such as healthcare, finance, and autonomous driving, information is often distributed across multiple modalities or views [^1] The terms modality and view will be used interchangeably in this paper. To effectively leverage this diverse information, multimodal AI models have been developed (Li et al.,, 2024). However, deep neural networks are prone to making wrong decisions with very high confidence (Guo et al.,, 2017; Hein et al.,, 2019). In safety-critical areas, such as healthcare or autonomous driving, this kind of overconfidence can lead to catastrophic results. Hence, when deploying these models in safety-critical areas, it is essential to understand the true uncertainty of the predictions.
@@
To address these issues, we propose a novel uncertainty aware multimodal fusion approach: Discounting Belief Fusion (DBF), which utilizes conflictive discounting of unreliable modalities, is order-invariant and can effectively scale to multiple modalities. Additionally, the proposed approach assigns higher uncertainties to conflictive decisions, which improves the reliability assessment of the final decision. We provide an experimental validation of our proposed approach, and open-source the underling source code in a public repository [^2] https://github.com/bezirganyan/DBF_uncertainty.
@@ 2.1 Uncertainty Quantification in Deep Learning
Evidential Deep Learning (EDL) (Sensoy et al.,, 2018), a single-network deterministic method, has recently gained popularity due to its easy adaptation to existing networks and its low training and testing times (Gao et al.,, 2024). EDL replaces the categorical distribution of the last layers of neural networks with a Dirichlet distribution and uses subjective logic (Jøsang,, 2016) to optimize the network parameters. Thanks to it’s ease of adaptation to existing architectures, and computational efficiency, it was also adapted to multi-modal models, which will be described in the next subsection.
@@ 2.2 Uncertainty Quantification in Multimodal and Multi-view Learning
In these approaches, incorporating a new modality decreases uncertainty regardless of the conflict between modalities, while a decision based on conflicting view can be less reliable in practice. To alleviate this issue, Xu et al., (2024) propose Evidential Conflictive Multiview Learning (ECML), which introduces an average pooling operation (averaging belief fusion in subjective logic). This method reduces the combined uncertainty if the new view has lower uncertainty and increases it if the new view has higher uncertainty. For two views, the combined uncertainty is the harmonic mean of the individual view uncertainties. However, it is possible to have two conflicting views with low uncertainty, resulting in low combined uncertainty when higher uncertainty would be expected. Additionally, the proposed fusion operation is not associative (Jøsang,, 2016), making the result of fusing more than two modalities highly dependent on the order of fusion. Huang et al., (2025) proposed a learnable discounting factor for modality reliability, but it uses a fixed value per modality and class, making it less adaptable to discrepancies between training and deployment. In contrast, our method computes reliability per sample, dynamically adjusting to misalignments and noise that may arise in real-world settings. Similar to these works, we also base our approach on evidential deep learning methods due to their computational efficiency and ease of adaptation. Nevertheless, we address shortcomings of the current approaches by presenting an order-invariant and scalable approach that effectively distinguishes between conflictive and non conflictive samples based on uncertainty estimates.
@@ 3.1 Subjective Logic and Uncertainty
Evidential Deep Learning uses Subjective Logic to model classification uncertainty by learning evidence scores for each class. Subjective Logic (Jøsang,, 2016) is the extension of Dempster–Shafer’s theory of belief functions (Dempster,, 1968; Shafer,, 1976) to allow a bijective mapping between a subjective opinion and the parameters of Dirichlet distribution. Specifically, for a classification problem of $K$ classes, the subjective opinion $\bm{\omega}$ is defined as a triplet $\bm{\omega}=({\bm{b}},u,{\bm{a}})$, where the ${\bm{b}}=({b}_{1},\dots,{b}_{K})^{T}$ represents the belief mass for each class, $u$ represents the uncertainty mass and ${\bm{a}}=({a}_{1},\dots,{a}_{K})^{T}$ represents the base rate or prior probability distribution over the classes. In scenarios where no prior knowledge about the classes is available (as is typical in most deep learning applications), a uniform base rate can be assumed. All these terms are strictly non negative and are less than equal to 1. Unlike probability theory, the belief masses do not sum up to one, however the additivity property dictates that $\sum_{i=1}^{K}{b}_{i}+u=1$. The base rates can usually be set to $1/K$ in case of absence of a prior belief the classes. The projected probability distribution for class $k$ can be obtained with $P_{k}={b}_{k}+{a}_{k}u,\forall k\in[1,\dots,K]$.
@@ 3.2 Generalized belief averaging
Let the dataset be defined as ${(x_{n}^{v},y_{n})}_{n=1}^{N},,v\in[1,\dots,V]$, where $x_{n}^{v}\in X$ represents the $n$-th data sample from the $v$-th modality, $y_{n}\in[0,1]^{K}$ is the corresponding one-hot encoded label vector, $V$ is the number of modalities, $N$ is the total number of samples, and $K$ is the number of classes. Having $V$ neural networks $f^{v}(x^{v})$ for each view, we obtain the view specific multinomial opinions $\bm{\omega}^{v}=f^{v}(x^{v})$, and the goal is to fuse these opinions into one final opinion: $\bm{\omega}^{\diamond V}=\bm{\omega}^{1}\diamond\dots\diamond\bm{\omega}^{V}$.
@@
This significantly reduces the scalability of the approach, since it means that $\bm{\omega}^{1\diamond 2\diamond 3}\neq\bm{\omega}^{3\diamond 2\diamond 1}$, which is counter-intuitive. To make the fusion process scalable and efficient for a many sources, we suggest to use the generalized version of the belief averaging fusion for multiple sources (Jøsang et al.,, 2017), with:
@@ 3.3 Discounted Belief Fusion
In this subsection, we demonstrate that the fusion approach described earlier does not adequately address the increased uncertainty in conflicting samples and propose the Discounted Belief Fusion (DBF) solution to mitigate this issue.
@@
The uncertainty estimate $u^{\diamond V}$ in equation 6 is the harmonic means of the individual uncertainties $u^{1},u^{2},\dots,u^{V}$.
@@
Looking back to the example in Figure 1, we can see that by using equation 13 to compute the discounting factor, the opinions that are in conflict with many other views (opinion 3 in the example), will be discounted much more than the other opinions.
Figure 1: The Conflict Matrix ${\bm{C}}$ for 4 opinions.
We can compute the updated belief masses and uncertainties with equation 13, and fuse the results with generalized belief averaging (equation 6).
@@
3.4 Training Multimoal Evidential Neural Networks
To fuse information from different modalities as described in the previous subsections, we first need to extract subjective opinions from each modality. For that, we construct evidential neural networks for each modality, which output non-negative evidence values to be later combined using a chosen fusion function.
@@ 3.4 Training Multimoal Evidential Neural Networks
where $\tilde{\alpha}_{n}=\mathbf{y}_{n}+\left(\mathbf{1}-\mathbf{y}_{n}\right)\odot\bm{\alpha}_{n}$ are the Dirichlet parameters after removing the non-misleading evidence, and $\Gamma(\cdot)$ is the gamma function. We can gradually increase the effect of the KL-Divergence loss with an annealing coefficient, to allow the network to explore the parameter, and avoid early convergence of the miscalssified samples to the uniform distribution (Sensoy et al.,, 2018). The annealing coefficient $\sigma_{t}$ can be chosen as $\sigma_{t}=\min(1.0,t/T)\in[0,1]$, where $t$ is the current epoch and $T$ is the annealing step.
@@ 4.1 Experimental setup
Five multimodal / multi-view datasets will be used for comparison: HandWritten (Duin,, 1998), CUB (Wah et al.,, 2011), Scene15 (Ali and Zafar,, 2018), Caltech101 (Li et al.,, 2022) and PIE (Gross et al.,, 2008). Please check the supplementary material for the descriptions and details about the datasets.
As baselines we use three evidential multi-view methods with different opinion aggregation strategies. The Trusted Multiview Classification (Han et al.,, 2021) uses Dempster’s combination rule or Belief Constraint Fusion (BCF), Trusted Multiview Deep Learning with Opinion Aggregation (Liu et al.,, 2022) uses Cumulative Belief Fusion (CBF), Evidential Conflictive Multiview Learning (Xu et al.,, 2024) which uses the Belief Averaging Fusion (BAF) and the Generalized Belief Averaging Fusion (GBAF, desribed in subsection 3.2).
@@ 4.2 Experimental Results
In the ablation studies, we tested all models using both softplus (for consistency with prior work) and exponential activation functions (see supplementary material for details). Results show that while the exponential activation improves model performance, it alone does not effectively separate uncertainties between conflictive and non-conflictive samples.
Our experiments showed that the proposed approach does not impact classification accuracy (see supplementary material). In most cases, the fusion function had minimal effect on final accuracy, allowing the proposed method to deliver improved uncertainty estimates without compromising the predictive performance of the models.
@@ Appendix B Training Details and Hyperparameters
We utilize 2-layer fully connected neural networks with ReLU activations for retrieving the evidences for each modality. In the final layer, we use either a softplus activation function (for baseline models, as proposed in their respective papers) or an exponential activation, capped at $10^{13}$ for numerical stability. To maintain differentiability, the activation function is modified to $f(x)=\frac{10^{13}}{1+10^{13}e^{-x}}$, which behaves exponentially for small inputs but gradually flattens as $x$ increases. To determine the optimal hyperparameters, we conducted a hyperparameter search using 10-fold cross-validation. To minimize time and resource consumption, this search was only performed on the ECML model Xu et al., (2024). This ensures that any slight performance differences due to hyperparameters would favor the baseline rather than our proposed approach. The selected hyperparameters are presented in Table 3. All experiments have been preformed 10 times with different random seeds, and the average and the standard deviation values have been reported in the results.
@@
Appendix C Details about the dataset

### v2

@@ Abstract
Multimodal AI models are increasingly used in fields like healthcare, finance, and autonomous driving, where information is drawn from multiple sources or modalities such as images, texts, audios, videos. However, effectively managing uncertainty—arising from noise, insufficient evidence, or conflicts between modalities—is crucial for reliable decision-making. Current uncertainty-aware machine learning methods leveraging, for example, evidence averaging, or evidence accumulation underestimate uncertainties in high-conflict scenarios. Moreover, the state-of-the-art evidence averaging strategy is not order invariant and fails to scale to multiple modalities. To address these challenges, we propose a novel multimodal learning method with order-invariant evidence fusion and introduce a conflict-based discounting mechanism that reallocates uncertain mass when unreliable modalities are detected. We provide both theoretical analysis and experimental validation, demonstrating that unlike the previous work, the proposed approach effectively distinguishes between conflicting and non-conflicting samples based on the provided uncertainty estimates, and outperforms the previous models in uncertainty-based conflict detection.
@@ 1 INTRODUCTION
Recently, the use of Artificial Intelligence (AI) has surged dramatically for both autonomous and assisted decision-making. In fields such as healthcare, finance, and autonomous driving, information is often distributed across multiple modalities or views [^1] The terms modality and view will be used interchangeably in this paper. To effectively leverage this diverse information, multimodal AI models have been developed (Li et al.,, 2024; Xiao et al.,, 2020; Lee and Yoo,, 2020; Chergui et al.,, 2024; Krones et al.,, 2025). However, deep neural networks are prone to making wrong decisions with very high confidence (Guo et al.,, 2017; Hein et al.,, 2019). In safety-critical areas, such as healthcare or autonomous driving, this kind of overconfidence can lead to catastrophic results. Hence, when deploying these models in safety-critical areas, it is essential to understand the true uncertainty of the predictions.
@@
To address these issues, we propose a novel uncertainty aware multimodal fusion approach: Discounting Belief Fusion (DBF), which utilizes conflictive discounting of unreliable modalities, is order-invariant and can effectively scale to multiple modalities. Additionally, the proposed approach assigns higher uncertainties to conflictive decisions, which improves the reliability assessment of the final decision. We provide an experimental validation of our proposed approach, and open-source the source code in a public repository [^2] https://github.com/bezirganyan/DBF_uncertainty.
@@ 2.1 Uncertainty Quantification in Deep Learning
Evidential Deep Learning (EDL) (Sensoy et al.,, 2018), a single-network deterministic method, has recently gained popularity due to its easy adaptation to existing networks and its low training and testing times (Gao et al.,, 2024). EDL replaces the categorical distribution of the last layers of neural networks with a Dirichlet distribution and uses subjective logic (Jøsang,, 2016) to optimize the network parameters. Thanks to its ease of adaptation to existing architectures, and computational efficiency, it was also adapted to multi-modal models, which will be described in the next subsection.
@@ 2.2 Uncertainty Quantification in Multimodal and Multi-view Learning
In these approaches, incorporating a new modality decreases uncertainty regardless of the conflict between modalities, while a decision based on conflicting view can be less reliable in practice. To alleviate this issue, Xu et al., (2024) propose Evidential Conflictive Multiview Learning (ECML), which introduces an average pooling operation (averaging belief fusion in subjective logic). This method reduces the combined uncertainty if the new view has lower uncertainty and increases it if the new view has higher uncertainty. For two views, the combined uncertainty is the harmonic mean of the individual view uncertainties. However, it is possible to have two conflicting views with low uncertainty, resulting in low combined uncertainty when higher uncertainty would be expected. Additionally, the proposed fusion operation is not associative (Jøsang,, 2016), making the result of fusing more than two modalities highly dependent on the order of fusion. Huang et al., (2025) proposed a learnable discounting factor for modality reliability, but it uses a fixed value per modality and class, making it less adaptable to discrepancies between training and deployment. In contrast, our method computes reliability per sample, dynamically adjusting to misalignments and noise that may arise in real-world settings. Similar to these works, we also base our approach on evidential deep learning methods due to their computational efficiency and ease of adaptation. Nevertheless, we address shortcomings of the current approaches by presenting an order-invariant and scalable approach that effectively distinguishes between conflictive and non-conflictive samples based on uncertainty estimates.
@@ 3.1 Subjective Logic and Uncertainty
Evidential Deep Learning uses Subjective Logic to model classification uncertainty by learning evidence scores for each class. Subjective Logic (Jøsang,, 2016) is the extension of Dempster–Shafer’s theory of belief functions (Dempster,, 1968; Shafer,, 1976) to allow a bijective mapping between a subjective opinion and the parameters of Dirichlet distribution. Specifically, for a classification problem of $K$ classes, the subjective opinion $\bm{\omega}$ is defined as a triplet $\bm{\omega}=({\bm{b}},u,{\bm{a}})$, where the ${\bm{b}}=({b}_{1},\dots,{b}_{K})^{T}$ represents the belief mass for each class, $u$ represents the uncertainty mass and ${\bm{a}}=({a}_{1},\dots,{a}_{K})^{T}$ represents the base rate or prior probability distribution over the classes. In scenarios where no prior knowledge about the classes is available (as is typical in most deep learning applications), a uniform base rate can be assumed. All these terms are strictly non-negative and are less than or equal to one. Unlike probability theory, the belief masses do not sum up to one, however the additivity property dictates that $\sum_{i=1}^{K}{b}_{i}+u=1$. The base rates can usually be set to $1/K$ in case of absence of a prior belief on the classes. The projected probability distribution for class $k$ can be obtained with $P_{k}={b}_{k}+{a}_{k}u,\forall k\in[1,\dots,K]$.
@@ 3.2 Generalized Belief Averaging
Let the dataset be defined as ${(x_{n}^{v},y_{n})}_{n=1}^{N},\ \ v\in[1,\dots,V]$, where $x_{n}^{v}\in X$ represents the $n$-th data sample from the $v$-th modality, $y_{n}\in[0,1]^{K}$ is the corresponding one-hot encoded label vector, $V$ is the number of modalities, $N$ is the total number of samples, and $K$ is the number of classes. Having $V$ neural networks $f^{v}(x^{v})$ for each view, we obtain the view-specific multinomial opinions $\bm{\omega}^{v}=f^{v}(x^{v})$, and the goal is to fuse these opinions into one final opinion: $\bm{\omega}^{\diamond V}=\bm{\omega}^{1}\diamond\dots\diamond\bm{\omega}^{V}$.
@@
As an example of Proposition 1, consider a scenario with three modalities. For a specific class $k$, the first modality provides an evidence value of 3, the second modality an evidence of 5, and the third modality an evidence of 10. When fusing these evidences using BAF, we compute:
$e_{k}^{1\diamond 2\diamond 3}=\frac{\left(\frac{3+5}{2}\right)+10}{2}=\frac{3+5}{4}+\frac{10}{2}=7$
Here, the evidence values from the first two modalities (3 and 5) are divided by 4, while the evidence value from the third modality (10) is divided by 2. Consequently, the contributions from the first two modalities are weighted only half as much as the third modality. This imbalance leads to a lack of commutativity in the fusion process, meaning that: $e_{k}^{1\diamond 2\diamond 3}\neq e_{k}^{3\diamond 2\diamond 1}$. To illustrate, if we change the fusion order:
$e_{k}^{3\diamond 2\diamond 1}=\frac{\left(\frac{10+5}{2}\right)+3}{2}=5.25$
This discrepancy is counter-intuitive and significantly hinders the scalability of the algorithm.
To make the fusion process scalable and efficient for many sources, we suggest to use the generalized version of the belief averaging fusion for multiple sources (Jøsang et al.,, 2017), with:
@@
Figure 1: The general pipeline of the Discounted Belief Fusion: First, subjective opinions are being formed for each modality using unimodal classifiers. Then conflict matrix is computed, which is then used to compute discounting factors for each modality. Finally, the beliefs of each modality are being adjusted using the discounting factors and are fused together using generalized belief averaging.
@@ 3.3 Discounted Belief Fusion
In this subsection, we demonstrate that the fusion approach described earlier does not adequately address the increased uncertainty in conflicting samples and propose the Discounted Belief Fusion (DBF) solution to mitigate this issue. The general pipeline of the DBF approach is shown in Figure 1.
@@
The uncertainty estimate $u^{\diamond V}$ in equation 6 is the harmonic mean of the individual uncertainties $u^{1},u^{2},\dots,u^{V}$.
@@
Looking back to the example in Figure 2, we can see that by using equation 13 to compute the discounting factor, the opinions that are in conflict with many other views (opinion 3 in the example), will be discounted much more than the other opinions. We can compute the updated belief masses and uncertainties with equation 13, and fuse the results with generalized belief averaging (equation 6).
@@
3.4 Training Multimodal Evidential Neural Networks
To fuse information from different modalities as described in the previous subsections, we first need to extract subjective opinions from each modality. For that, we construct evidential neural networks for each modality, which output non-negative evidence values to be later combined using a chosen fusion function. Simultaneous availability of the modalities during training is preferred but not mandatory. If simultaneous availability is not possible, one can effectively integrate new modalities without complete retraining, by training only a classifier for the new modality, and if necessary, fine-tuning the other modalities.
@@ 3.4 Training Multimodal Evidential Neural Networks
where $\tilde{\alpha}_{n}=\mathbf{y}_{n}+\left(\mathbf{1}-\mathbf{y}_{n}\right)\odot\bm{\alpha}_{n}$ are the Dirichlet parameters after removing the non-misleading evidence, and $\Gamma(\cdot)$ is the gamma function. We can gradually increase the effect of the KL Divergence loss with an annealing coefficient, to allow the network to explore the parameter, and avoid early convergence of the misclassified samples to the uniform distribution (Sensoy et al.,, 2018). The annealing coefficient $\sigma_{t}$ can be chosen as $\sigma_{t}=\min(1.0,t/T)\in[0,1]$, where $t$ is the current epoch and $T$ is the annealing step.
@@ 4.1 Experimental setup
Five multimodal / multi-view datasets will be used for comparison: HandWritten (Duin,, 1998), CUB (Wah et al.,, 2011), Scene15 (Ali and Zafar,, 2018), Caltech101 (Li et al.,, 2022) and PIE (Gross et al.,, 2008). Please check the supplementary material for the descriptions and details about the datasets. As baselines we use three evidential multi-view methods with different opinion aggregation strategies. The Trusted Multiview Classification (Han et al.,, 2021) uses Dempster’s combination rule or Belief Constraint Fusion (BCF), Trusted Multiview Deep Learning with Opinion Aggregation (Liu et al.,, 2022) uses Cumulative Belief Fusion (CBF), Evidential Conflictive Multiview Learning (Xu et al.,, 2024) uses the Belief Averaging Fusion (BAF) and the Generalized Belief Averaging Fusion (GBAF, described in subsection 3.2).
@@ 4.2 Experimental Results
In the ablation studies, we tested all models using both softplus (for consistency with prior work) and exponential activation functions (see Appendix E for details). Results show that while the exponential activation improves model performance, it alone does not effectively separate uncertainties between conflictive and non-conflictive samples.
Our experiments showed that the proposed approach does not impact classification accuracy (see Appendix D). In most cases, the fusion function had minimal effect on final accuracy, allowing the proposed method to deliver improved uncertainty estimates without compromising the predictive performance of the models.
Table 3: Training times of different algorithms for each dataset. The means and standard deviations of 10 runs are reported in seconds.
Dataset BAF BCF DBF (Our)
Scene15 $757\pm 22$ $854\pm 25$ $798\pm 11$
CalTech $413\pm 2$ $419\pm 5$ $455\pm 3$
PIE $133\pm 2$ $146\pm 2$ $140\pm 2$
HandWritten $607\pm 4$ $752\pm 6$ $602\pm 12$
CUB $94\pm 0.5$ $99\pm 1$ $105\pm 1$
As discussed in the Appendix A, the conflict matrix computation has a quadratic time complexity relative to the number of modalities. However, since the number of modalities is typically not large, the practical time overhead of using DBF remains low. To validate this, we present the training times of various baseline algorithms alongside the DBF method in Table 3. As observed, DBF is not consistently slower than other algorithms, and even when it is, the difference is not much.
@@ 5 CONCLUSION
In future work, we plan to further validate our approach using larger benchmark datasets designed for uncertainty-aware multimodal learning, such as the LUMA dataset Bezirganyan et al., (2024). Additionally, we plan to explore modality interactions and dependency structures, which could further enhance our uncertainty quantification strategy.
Acknowledgements
Centre de Calcul Intensif d’Aix-Marseille is acknowledged for granting access to its high-performance computing resources.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. Yes
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. Yes. Please refer to the Appendix A.
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Yes
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. Yes
    2. (b)
      Complete proofs of all theoretical results. Yes
    3. (c)
      Clear explanations of any assumptions. Yes
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). Yes, code included in an anonymized repository
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). Yes. Please refer to Appednix B.
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). Yes
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). Yes
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. Yes
    2. (b)
      The license information of the assets, if applicable. Yes
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. Not Applicable
    4. (d)
      Information about consent from data providers/curators. Not Applicable
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not Applicable
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. Not Applicable
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not Applicable
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not Applicable
@@ Appendix B TRAINING DETAILS AND HYPERPARAMETERS
We utilize 2-layer fully connected neural networks with ReLU activations for retrieving the evidences for each modality. In the final layer, we use either a softplus activation function (for baseline models, as proposed in their respective papers) or an exponential activation, capped at $10^{13}$ for numerical stability. To maintain differentiability, the activation function is modified to $f(x)=\frac{10^{13}}{1+10^{13}e^{-x}}$, which behaves exponentially for small inputs but gradually flattens as $x$ increases. To determine the optimal hyperparameters, we conducted a hyperparameter search using 10-fold cross-validation. To minimize time and resource consumption, this search was only performed on the ECML model Xu et al., (2024). This ensures that any slight performance differences due to hyperparameters would favor the baseline rather than our proposed approach. The selected hyperparameters are presented in Table 4. All experiments have been performed 10 times with different random seeds, and the average and the standard deviation values have been reported in the results.
@@
Appendix C DETAILS ABOUT THE DATASETS